# Federated Causal Inference from Multi-Site Observational Data via Propensity Score Aggregation

### v1

@@ 4.1 Oracle federated estimators
As discussed in the previous section, existing federated causal inference methods often rely on restrictive assumptions—such as a common propensity score across sites (Guo et al.,, 2025), site differences limited to intercept shifts (Yin et al.,, 2025), or predefined shared structures (Xiong et al.,, 2023). In practice, treatment assignment often varies across sites due to differences in norms, resources, or clinical practices. To accommodate this heterogeneity, we express the global propensity score as a weighted combination of site-specific scores, capturing local treatment policies.
Specifically, the global propensity score can be expressed as a weighted combination of local scores using two alternative weighting schemes (see Appendix A.4). The first relies on Membership Weights (MW), representing the probability of belonging to each site given the covariates:
$e(X)=\sum_{k=1}^{K}\underbrace{\mathbb{P}(H=k\mid X)}_{\text{membership weights}}e_{k}(X).$ (3)
The second uses Density ratio Weights (DW), based on the ratio between the covariate density at site $k$, denoted $f_{k}$, and the overall population density $f$:
$e(X)=\sum_{k=1}^{K}\underbrace{\rho_{k}\frac{f_{k}(X)}{f(X)}}_{\text{density weights}}e_{k}(X).$ (4)
Building on Eq. 3 and 4, we define our oracle Federated (A)IPW estimators, denoted Fed-(A)IPW, where the corresponding weights will be estimated via federated learning (see Section 4.2).
Definition 3 (Oracle federated estimators).
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{IPW}},\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{AIPW}},$ (5)
@@
where $\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{IPW}}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e)$, and $\hat{\tau}^{\mathrm{fed}(k)}_{\mathrm{AIPW}}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})$ rely on the global propensity score $e(X)=\sum_{k=1}^{K}\omega_{k}(X)e_{k}(X)$ with federated weights $\{\omega_{k}(X)\}_{k}$ that are either the membership (Eq. 3) or the density ratio (Eq. 4) weights.
Crucially, our decompositions of the global propensity score in Eq. 3 and 4 enable a flexible combination of locally estimated propensity scores $e_{k}$ and globally learned federation weights $\omega_{k}(x)$. Theorem 3 (proved in Appendix A.5) establishes that this approach, in the oracle setting, matches the efficiency of centralized estimators.
Theorem 3 (Equality of centralized and federated (A)IPW estimators).
@@
Under Assumptions 1, 2 and 3, the oracle federated estimators (Def. 3) are equal to the oracle centralized estimators (Def. 1).
@@
We further show in Appendix A.6 that even when local overlap (Assumption 4) holds, our federated estimators achieve lower variance than meta-analysis estimators.
Theorem 4 (Variance comparison between meta-analysis and federated estimators).
Under Assumptions 1, 2 and 4, both federated and meta-analysis estimators can be computed, and
$\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{*}]=\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}\right],$
$\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{*}]=\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}\right],$
@@
with equality when the local propensity scores are equal.
We provide two main reasons for this improvement. First, our decomposition of $e$ in Eq.3 marginalizes over $H\mid X$, effectively removing the need to adjust for site membership $H$. This reduces variance by avoiding unnecessary conditioning. Second, our federated approach inherently yields better overlap than the siloed meta-analysis framework, as formalized in the following theorem.
@@
Theorem 5 (proved in Appendix A.7) establishes that global overlap is always at least as good as the worst local overlap. This implies that even when local overlap holds, sites with poor overlap still benefit from the federated approach: the global propensity scores $e(X)$ are more bounded away from 0 and 1 than the local scores $\{e_{k}(X)\}_{k\in[K]}$. Notably, sites with poor individual overlap can even improve the overall overlap of the federated population, as illustrated in the following example.
Example 1. Let $K=2$ with $X_{i}=1$ in both sites, $n_{1}=n_{2}$, $\mathbb{P}(H_{i}\mid X_{i})=0.5$, $e_{1}(X_{i})=0.99\times X_{i}$ and $e_{2}(X_{i})=0.01\times X_{i}$, leading to $e(X_{i})=0.5\times e_{1}(X_{i})+0.5\times e_{2}(X_{i})=0.5\times X_{i}$. Then, $\mathcal{O}_{1}=\mathcal{O}_{2}=\left(0.99\times 0.01\right)^{-1}\approx 101$, while $\mathcal{O}_{\mathrm{global}}=\left(0.5\times 0.5\right)^{-1}=4$. This illustrates how heterogeneity among local propensity scores over a shared population can enhance global overlap.
@@ 4.2 Federated estimation
We now move beyond oracle estimators and describe how to implement the Fed-(A)IPW estimators in a practical federated learning setting. Our method follows a two-step procedure (which can be executed in parallel): each site $k$ first estimates and shares a local propensity score function $\hat{e}_{k}(x)$; then, global scores $\hat{e}(x)$ are computed as weighted averages of the local scores, using weights $\{\hat{\omega}_{k}(x)\}_{k\in[K]}$ estimated in a federated manner. For Fed-AIPW, a third step is needed to train the outcome models $\hat{\mu}_{0},\hat{\mu}_{1}$ with federated learning. Details on how to estimate $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$, along with $\hat{\mu}_{0},\hat{\mu}_{1}$, are provided in the following subsections.
4.2.1 Estimation of the local propensity scores
Local propensity scores can be estimated using any probabilistic binary classification method. Common choices include logistic regression for parametric estimation and generalized random forests for non-parametric estimation (Lee et al.,, 2010). A key advantage of our approach is its flexibility: different estimation methods can be used across sites, tailored to local data characteristics or computational constraints.
As previously discussed, this procedure does not require the overlap assumption to hold locally. That is, local propensity scores can legitimately approach 0 or 1—as long as the global score remains sufficiently bounded away from these extremes. This property makes the approach suitable for scenarios involving external control arms (Center for Drug Evaluation and Research and Center for Biologics Evaluation and Research and Oncology Center of Excellence,, 2023; European Medicines Agency,, 2023) where some sites have $e_{k}(X_{i})=0$ for all $i$ in the control group, yet still contribute to the global analysis.
4.2.2 Estimation of the federated weights
Membership weights.
We begin with membership weights $\{\omega_{k}(x)=\mathbb{P}(H_{i}=k\mid X=x)\}_{k}$, which represent the probability of site membership given covariates. These weights can be estimated without access to individual-level data using any federated algorithm for multi-class classification models that can output probabilities.
Federated versions of parametric models—such as logistic regression or deep neural networks—trained via gradient-based empirical risk minimization are well-suited for this task and readily supported in practice (Kairouz et al.,, 2021). Additionally, non-parametric models like random forests (Hauschild et al.,, 2022) and gradient-boosted trees (Li et al.,, 2020) have been adapted to the federated setting, offering more flexibility in capturing complex decision boundaries.
In our experiments, we use a simple multinomial logistic regression model trained via the standard Federated Averaging (FedAvg) algorithm (McMahan et al.,, 2017). This approach is easy to deploy in a server–client architecture, remains interpretable, and requires exchanging only $TKd$ floats, where $T$ is the number of communication rounds. We refer to Appendix B for details.
Density ratios weights.
We now turn to density ratio weights $\{\omega_{k}(x)=\rho_{k}\frac{f_{k}(x)}{f(x)}\}_{k}$, where $f_{k}$ and $f$ denote the local and global covariate densities, respectively. Although various methods exist for estimating density ratios (Sugiyama et al.,, 2012), we are not aware of any that are directly compatible with federated settings. Standard non-parametric techniques, such as kernel density estimation, typically require sharing raw data, which violates the decentralized constraints we impose.
We thus propose to estimate each local density $f_{k}$ using a parametric model (e.g. via maximum likelihood estimation). In our experiments, we assume that covariates are normally distributed within each site. Each site $k$ locally estimates its mean $\hat{\mu}_{k}=1/n_{k}\sum_{i=1}^{n_{k}}X_{i}$ and covariance matrix $\hat{\Sigma}_{k}=1/n_{k}\sum_{i=1}^{n_{k}}(X_{i}-\hat{\mu}_{k})(X_{i}-\hat{\mu}_{k})^{\top}$, then transmits these to the central server. Noting that the global density $f$ is a mixture of the $K$ local densities weighted by site proportions $\rho_{k}$ (which can be estimated by $n_{k}/n$), we estimate the density ratios weights as $\hat{\omega}_{k}(x)=\frac{\frac{n_{k}}{n}\hat{f}_{k}(x)}{\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{f}_{k}(x)}$, where each local density is given by
$\hat{f}_{k}(x)$ $=\frac{1}{(2\pi)^{d/2}|\hat{\Sigma}_{k}|^{1/2}}\times\exp\Big{(}-\frac{1}{2}(x-\hat{\mu}_{k})^{\top}\hat{\Sigma}_{k}^{-1}(x-\hat{\mu}_{k})\Big{)}.$
In contrast to the estimation of membership weights, this approach requires only a single round of communication, but involves sending $Kd+Kd^{2}$ floats, which can be significant for high-dimensional data.
More broadly, we recommend using a parametric approach to density ratio estimation when prior knowledge about the local distributions is available and well-justified. However, in the absence of reliable distributional assumptions, membership weights often offer a more robust and practical alternative. They allow for flexible parameterizations and can be estimated through classification models, which are generally easier to learn than densities—particularly in high-dimensional settings where density estimation suffers from high sample complexity and poor scalability (Sugiyama et al.,, 2012).
4.2.3 Estimation of Fed-AIPW
Fed-AIPW requires estimating the outcome models $\mu_{1}$ and $\mu_{0}$ in addition to the propensity components $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$. The estimation procedure for $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$ follows the same steps outlined in Sections 4.2.1 and 4.2.2. For the outcome models, one can estimate $\mu_{0}$ and $\mu_{1}$ globally using federated gradient descent, as in Khellaf et al., (2025), with parametric models such as logistic regression for binary outcomes or linear regression for continuous outcomes. As with the estimation of membership weights, our framework also supports more flexible outcome modeling using a broad range of function classes, including neural networks, decision forests, and other gradient-based models adapted to the federated setting.
@@ 5 Experiments
We consider $K=3$ sites and $d=10$ covariates. Two distinct data-generating processes (DGPs) are used in the simulations. In DGP A, each site $k$ independently samples $n_{k}=500$ individuals from a site-specific multivariate Gaussian distribution $\mathcal{N}(\mu_{k},\Sigma_{k})$. In DGP B, a total of $n=1500$ individuals are first drawn from a bimodal Gaussian mixture and then assigned to sites using a multinomial logistic model based on their covariates. We vary the degree of overlap within sites to reflect different practical scenarios: No local overlap (the second site has no treated individuals, hence $\mathcal{O}_{2}=\infty$), Poor local overlap ($\mathcal{O}_{2}\approx 10^{7}$), and Good local overlap ($\mathcal{O}_{2}\approx 4.6$). The outcome models $\mu_{1}$ and $\mu_{0}$ are shared across sites and specified as linear functions with interaction terms. For comparison, we also generate data consistent with the setting described in Xiong et al., (2023) (Figure 5). All reported results are aggregated over 1,500 simulation runs. Complete details of the simulation setup are provided in Appendix C.
We evaluate several estimators: (i) our proposed Fed-IPW and Fed-AIPW, with either Membership Weights (MW) (Def. 3) estimated via federated multinomial logistic regression (well-specified in DGP B), or Density Ratio Weights (DW) (Def. 4) based on Gaussian models (well-specified in DGP A); (ii) the Centralized Oracle (Def. 1); (iii) meta-analysis IPW and AIPW estimators with sample size weighting (Meta-SW) (Def. 2); and (iv) the one-shot inverse variance weighted IPW estimator (1S-IVW) from Xiong et al., (2023), evaluated under a favorable setting for this method—specifically, we assume that certain propensity score parameters are shared across sites. For all estimators, propensity scores are estimated with logistic regression, and outcome models with linear regression.
The results for the three degrees of local overlap are shown in Figures 5, 5 and 5, with the top (respectively bottom) row showing DGP A (respectively DGP B). Before discussing each setting separately, we start by some general observations. First, Fed-IPW-DW is unbiased in DGP A but biased in DGP B, while our Fed-IPW-MW is unbiased in DGP B but biased in DGP A, as expected. Note that Fed-IPW-MW could be made unbiased by using more flexible models (e.g., neural networks) to estimate membership probabilities. Second, our Fed-AIPW estimators remain unbiased across all overlap levels thanks to their double robustness property: despite misspecified linear outcome models, well-estimated propensity scores ensure consistency. Finally, in most cases our Fed-IPW estimators exhibit lower variance than the oracle IPW estimator. This is because they leverage estimated global propensity scores, a strategy known to reduce variance in IPW estimation (Hirano et al.,, 2003).
In the No local overlap setting (Figure 5), meta-analysis estimators are undefined due to the absence of treated individuals at one site. In contrast, our federated estimators remain unbiased under both DGPs, as Assumption 3 holds (with global overlap $\mathcal{O}_{\mathrm{global}}\approx 6.22$). In the Poor local overlap setting (Figure 5), site 2 lacks sufficient overlap, leading to bias and instability in the meta-analysis estimators, including Meta-AIPW as both the propensity weights and local outcome models are inaccurate. However, this issue is mitigated in the global dataset (see Figure 5 in the Appendix), allowing our federated estimators to perform reliably. Finally, in the Good local overlap setting (Figure 5), while all competitors are unbiased, our federated methods exhibit lower variance.
In Figure 5, we consider a specific setting in which all local propensity score functions share a common subset of 5 out of 10 logistic regression coefficients with data generated from DGP B. This setup aligns with the assumptions of the 1S-IVW method proposed by Xiong et al., (2023), which relies on prior knowledge of the shared coefficients—an assumption our method does not require. Nevertheless, Fed-MW is unbiased and exhibits the smallest variance.
Figure 2: Synthetic data: DGP A (top panel) and B (bottom).
@@
Figure 3: Comparison to Xiong et al., (2023)
Figure 4: Results of Traumabase experiment.
Real data.
We apply our methods to the multi-site observational Traumabase cohort, which contains data on patients with traumatic brain injury (Mayer et al.,, 2020; Colnet et al.,, 2024). We study the effect of tranexamic acid (TA), a drug used to reduce bleeding, on the binary outcome of mortality among trauma patients. For this analysis, we select the $K=4$ centres with highest number of treated patients (min. 47), totaling 472 treated and 5,531 control patients, each with 17 measured covariates. We standardize the covariates to have zero mean and unit variance in a federated manner by computing and sharing the means and variances of the covariates across sites.
We focus on AIPW estimators. For Fed-MW and Fed-DW, we estimate the $e_{k}$’s with local logistic regression, while the outcome models $\mu_{1},\mu_{0}$ and the membership weights are learned with a logistic regression model trained with FedAvg (Algorithm 1 in the Appendix) with $T=5000$ rounds, $E=1$ local steps and learning rate $\eta=0.1$. The density ratio weights are computed under a Gaussian model for the covariates. We compare our estimators to Meta-SW and centralized AIPW estimators using also linear and logistic models. To construct empirical confidence intervals, we perform 1000 bootstrap resamples. The results in Figure 4 demonstrate that our federated approach yields estimates comparable to those of the centralized method, while exhibiting lower variance than Meta-SW. Notably, prior analyses using the combined data from a larger set of sites have reported a non-significant average treatment effect (Mayer et al.,, 2020; Colnet et al.,, 2024), which is consistent with our findings on this data subset.

### v2

@@ 3.1 Oracle Federated Estimators
As discussed before, existing federated causal inference methods often rely on restrictive assumptions—such as a common propensity score across sites (Guo et al., 2025), site differences limited to intercept shifts (Yin et al., 2025), or predefined shared structures (Xiong et al., 2023). In practice, treatment assignment frequently varies across sites due to differences in norms, resources, or clinical practices. To properly account for this heterogeneity, the global propensity score must be expressed as a weighted combination of the site-specific scores (see Appendix A.4). A first choice of weights are the density ratio weights (DW) $\omega_{k}^{\mathrm{DW}}$:
$e(x)=\sum_{k=1}^{K}\underbrace{\rho_{k}\frac{f_{k}(x)}{f(x)}}_{=\omega_{k}^{\mathrm{DW}}(x)}e_{k}(x),$ (3)
where $f_{k}$ and $f$ are the covariate densities locally and globally. Similar weights are used in transportability methods that reweight data to match a target population (Han et al., 2023, 2025; Guo et al., 2024), though here the objective is to recover the global propensity score across the super-population defined by the $K$ participating sites. Unfortunately, DW estimation in a federated setting requires modeling the $f_{k}$’s, which entails strong distributional assumptions and becomes challenging in high dimensions.
Instead, we propose the Membership Weights (MW) $\omega_{k}^{\mathrm{MW}}$:
$e(x)=\sum_{k=1}^{K}\underbrace{\mathbb{P}(H=k\mid X=x)}_{\displaystyle=\omega_{k}^{\mathrm{MW}}(x)}e_{k}(x),$ (4)
which represent the probability of site membership given the covariates. Unlike DW, MW do not require explicit density modeling and can be estimated directly via federated parametric (e.g., logistic regression, neural networks) or non-parametric (e.g., gradient-boosted trees) classification models, providing a flexible, communication-efficient alternative and making it the preferred choice for federated settings. We refer to Section 3.2 for more details on the federated estimation of MW and its advantages over DW.
Equations 3 and 4 enable combining locally estimated propensity scores $e_{k}$ into a global propensity score using globally learned weights $\omega_{k}(x)$. Building on this decomposition, we define our oracle Federated IPW and AIPW estimators (Fed-(A)IPW). We define the Oracle federated estimators as follows:
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed}(k)},\qquad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed}(k)},$ (5)
@@
where $\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed}(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e)$ and $\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed}(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})$ rely on the global propensity score $e(X)=\sum_{k=1}^{K}\omega_{k}(X)e_{k}(X)$.
Theorem˜3 (proved in Section˜A.5) establishes that, in the oracle setting, Fed-(A)IPW estimators attain the same efficiency as their centralized counterparts.
Theorem 3.
@@
Under Assumptions 1, 2, and 3, the oracle federated estimators (Equation 5) are identical to the oracle centralized estimators (Equation 1).
@@
Theorem˜4 (proved in Appendix A.6) further shows that even when local overlap (Assumption 4) holds, federated estimators have lower variance than meta-analysis estimators.
Theorem 4.
Under Assumptions 1, 2 and 4, we have:
$\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}],\quad\quad\quad\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}],$
@@
with equality when the local propensity scores are identical across sites.
This variance reduction arises for two reasons. First, decomposing $e$ as a weighted sum of $e_{k}$’s marginalizes over $H\mid X$, eliminating unnecessary adjustment for site membership and thereby reducing variance. Second, our federated approach improves overlap compared with meta-analysis, as formalized below.
@@
Theorem˜5 (proved in Appendix A.7) shows that global overlap is always at least as good as the worst local overlap. Even when local overlap holds, sites with poor overlap benefit from the federated approach because the global score $e(x)$ is more bounded away from 0 and $1$ than the local scores $\{e_{k}(x)\}_{k\in[K]}$. Notably, sites with poor individual overlap can even improve the overall overlap of the federated population, as illustrated in the following example.
Example.
Let $K=2$ with $X_{i}=1$ in both sites, $\mathbb{P}(H_{i}\mid X_{i})=0.5$, $e_{1}(X_{i})=0.99$ and $e_{2}(X_{i})=0.01$, leading to $e(X_{i})=\sum_{k=1}^{2}0.5\times e_{k}(X_{i})=0.5$. Local overlaps are poor, $\mathcal{O}_{1}=\mathcal{O}_{2}=\left(0.99\times 0.01\right)^{-1}\approx 101$, whereas the global overlap is $\mathcal{O}_{\mathrm{global}}=\left(0.5\times 0.5\right)^{-1}=4$—the optimal value achieved in a randomized trial with $50\%$ treatment probability. This illustrates how heterogeneity in treatment assignments can enhance global overlap and enable more robust causal inference.
@@ 3.2 Federated estimation
We now move beyond oracle estimators and describe how to implement our Fed-(A)IPW estimators in a practical federated learning setting. Constructing the global score propensity score requires two steps, which can be executed in parallel: each site $k$ estimates and shares a local propensity score $\hat{e}_{k}(x)$; and the sites collaboratively estimate federated weights $\{\hat{\omega}_{k}(x)\}_{k\in[K]}$. Fed-AIPW adds a third step to train outcome models $\hat{\mu}_{0},\hat{\mu}_{1}$ via federated learning. We detail how to estimate $\{\hat{e}_{k}(x),\hat{\omega}_{k}(x)\}_{k}$ and $\hat{\mu}_{0},\hat{\mu}_{1}$ below.
Local propensity scores. Each $e_{k}$ can be estimated using any probabilistic binary classifier, either parametric (e.g., logistic regression or neural networks) or non-parametric ( (e.g., generalized random forests, Lee et al., 2010). A key advantage of our approach is flexibility: sites can use different estimation methods tailored to local data or computational constraints. It also does not require Assumption 4: local scores may approach 0 or 1 provided the global score remains bounded away from these extremes.This, in particular, enables the integration of external control arms (FDA, 2023; EMA, 2023), where some sites have $e_{k}(X)=0$ for all control patients yet still contribute to the global analysis.
Federated weights: density ratio vs. membership. Parametric density ratio weights $\omega_{k}^{\mathrm{DW}}(x)=\rho_{k}\frac{f_{k}(x)}{f(x)}$ can be implemented in a one-shot fashion by sharing local density parameters with the server, which then reconstructs the global mixture and computes the weights. A common choice is to assume parametric covariate distributions (say, Gaussian), estimate $(\hat{\mu}_{k},\hat{\Sigma}_{k})$ locally, and transmit them once to the server. This requires to communicate $O(Kd^{2})$ parameters and is highly sensitive to model misspecification—an issue that becomes critical in high dimensions. Nonparametric density estimation would relax these assumptions but is statistically inefficient and does not yet have practical federated implementations.
In contrast, our membership weights $\omega_{k}^{\mathrm{MW}}(x)=\mathbb{P}(H=k\mid X=x)$ can be learned with any probabilistic multiclass classifier trained federatively—for example, logistic regression for simplicity and interpretability, or neural networks to capture complex nonlinearities. Such models are readily supported by modern FL algorithms and software libraries. Using the standard FedAvg algorithm (McMahan et al., 2017) requires exchanging $TKP$ floats (training rounds $\times$ sites $\times$ model parameters), which is feasible for models of practical size and a large number of sites. Non-parametric classifiers like random forests (Hauschild et al., 2022) and gradient-boosted trees (Li et al., 2020) have also been adapted to the federated setting.
Outcome models. To construct the doubly robust Fed-AIPW estimator, $\mu_{0},\mu_{1}$ are trained federatively, as in Khellaf et al. (2025). As for MW, standard FL algorithms and librairies can be used to train a wide range of parametric and non-parametric supervised learning models.
Remark (Missing values).
Our approach can naturally accommodate missing data. Local propensity scores can be estimated consistently with logistic regression or random forests (Jiang et al., 2020; Josse et al., 2024). For membership weights, constant imputation combined with federated random forests provides a consistent solution (Le Morvan et al., 2021), whereas density ratio weights would require adapting more complex federated EM algorithms (Dieuleveut et al., 2021; Marfoq et al., 2021). Finally, doubly robust estimators with missing data can be obtained via non-parametric federated outcome models (Mayer et al., 2020).
@@ 4 Experiments
(d) Synthetic data: DGP A (top) and DGP B (bottom).
We consider $K=3$ sites and $d=10$ covariates. Two data-generating processes (DGPs) are used. In DGP A, each site $k$ independently samples $n_{k}=650$ individuals from a site-specific multivariate Gaussian distribution $\mathcal{N}(\mu_{k},\Sigma_{k})$. In DGP B, a total of $n=4000$ individuals are first drawn from a bimodal Gaussian mixture and then assigned to sites according to a multinomial logistic model based on their covariates. We vary within-site overlap to mimic different practical scenarios: No local overlap ($\mathcal{O}_{2}=+\infty$, the second site has no treated individuals), Poor local overlap ($\mathcal{O}_{2}\approx 10^{7}$), and Good local overlap ($\mathcal{O}_{2}\approx 4.6$). The outcome models $\mu_{1},\mu_{0}$ are shared across sites and specified as polynomial functions with interactions. For comparison, we also generate data consistent with the setting in Xiong et al. (2023) (Figure 2). All results are averaged over 800 simulation runs; full details are provided in Appendix C.
We evaluate our proposed Fed-IPW and Fed-AIPW using the MW weights estimated either via federated multinomial logistic regression (well specified in DGP B) or via a two-layer Neural Network (NN). We compare these methods against several competitors: Fed-IPW and Fed-AIPW with the alternative DW weights based on Gaussian density estimation (well specified in DGP A); the Centralized Oracle (Def. 1); meta-analysis IPW/AIPW with sample-size weighting (Meta-SW) (Def. 2); and the one-shot inverse-variance weighted IPW estimator (1S-IVW) of Xiong et al. (2023), evaluated under favorable conditions with shared propensity-score parameters across sites. For all estimators, propensity scores are fit via logistic regression and outcome models via linear regression, implying misspecification in the latter.
Figures 44 (top: DGP A; bottom: DGP B) summarize the results under the three overlap regimes. Before discussing each setting, we highlight several general observations. First, Fed-IPW-DW is unbiased only under DGP A, where the Gaussian specification holds, and exhibits bias under DGP B. By contrast, Fed-IPW-MW adapts across data-generating processes: with logistic membership weights it is unbiased in DGP B, and with a more flexible neural network classifier it is unbiased for both DGP A and DGP A, reflecting the greater modeling flexibility of MW relative to DW’s restrictive density specification. Second, Fed-AIPW enjoys its doubly robust property and remains unbiased across all overlap levels by double robustness; despite misspecified linear outcome models, accurate propensity estimation ensures consistency. Finally, Fed-IPW typically attains lower variance than the centralized oracle IPW, consistent with well-documented efficiency gains from using estimated propensity scores (Hirano et al., 2003).
In the No local overlap setting (Figure 4), meta-analysis estimators are undefined because one site has no treated individuals. Our federated estimators remain unbiased under both DGPs, as ˜3 holds (global overlap $\mathcal{O}_{\mathrm{global}}\approx 6.22$). In the Poor local overlap setting (Section˜4), site 2 exhibits weak overlap, leading to bias and instability in meta-analysis estimators, including Meta-AIPW, as both the propensity scores and local outcome models are inaccurate. This issue is mitigated in the global dataset (see Figure 4 in the Appendix), allowing our federated estimators to remain reliable. In the Good local overlap setting (Figure 4), all methods are unbiased, but federated estimators achieve the smallest variance.
Figure 2 considers a setting where all local propensity scores share a common subset of 5 of 10 logistic regression coefficients with data generated from DGP B. This setup matches the assumptions of the 1S-IVW method of Xiong et al. (2023), which relies on prior knowledge of the shared parameters to aggregate them—an assumption not required by our method. Fed-MW remains unbiased and attains the lowest variance, matching that of 1S-IVW, even without access to the additional information about shared parameters.
Additional simulations provided in Appendix E examine scenarios with more sites (Table 4), nonparametric estimation (Table 5), and local propensity model misspecification (Table 6), further confirming the robustness of our approach.
Figure 2: Comparison to Xiong et al. (2023) (IPW).
Figure 3: AIPW estimates on Traumabase.
Real data. We analyze the multi-site Traumabase cohort (Mayer et al., 2020; Colnet et al., 2024) to estimate the effect of tranexamic acid on mortality across $K=14$ centers. The local datasets are highly imbalanced in site sizes (106 to 2,092 patients) and treatment arms (e.g., site 11: 4 treated vs. 121 controls); there are 17 covariates and n=8,248n=8{,}248 patients in total, of whom 638 were treated. Covariates are standardized federatively by sharing site means and variances. We focus on AIPW estimators: local propensities $e_{k}$ are estimated via logistic regression; outcome models $\mu_{1},\mu_{0}$ are trained with FedAvg logistic regression (5,000 rounds, 1 local epoch, step size $\eta=0.1$); MW are learned with a FedAvg neural network (one hidden layer, 128 units), while DW are based on Gaussian density estimation. Competitors include Meta-SW (same local models, computed only on sites with enough treated units—where the number of treated observations exceeds the covariate dimension) and a centralized AIPW benchmark obtained using the R package grf’s probability_forest function (Tibshirani et al., 2018) on the pooled data. Empirical confidence intervals are constructed from 150 bootstrap resamples.
Figure 3 shows that our federated estimators closely match the centralized benchmark and exhibit lower variance than Meta-SW, which in this application departs from the centralized estimate. Among federated methods, Fed-MW is closest to the nonparametric centralized estimator—reflecting its more flexible modeling of propensities and membership—whereas Fed-DW is less stable, likely due to noisy per-site covariance estimates for a 17-dimensional Gaussian ($\approx 17^{2}$ parameters) in small sites.